### v1

# How Much are LLMs Contaminated?

Abstract
With the rise of Large Language Models (LLMs) in recent years, new opportunities are emerging, but also new challenges, and contamination is quickly becoming critical. Business applications and fundraising in AI have reached a scale at which a few percentage points gained on popular question-answering benchmarks could translate into dozens of millions of dollars, placing high pressure on model integrity. At the same time, it is becoming harder and harder to keep track of the data that LLMs have seen; if not impossible with closed-source models like GPT-4 and Claude-3 not divulging any information on the training set. As a result, contamination becomes a critical issue: LLMs’ performance may not be reliable anymore, as the high performance may be at least partly due to their previous exposure to the data. This limitation jeopardizes the entire progress in the field of NLP, yet, there remains a lack of methods on how to efficiently address contamination, or a clear consensus on prevention, mitigation and classification of contamination. In this paper, we survey all recent work on contamination with LLMs, and help the community track contamination levels of LLMs by releasing an open-source Python library named LLMSanitize implementing major contamination detection algorithms, which link is: https://github.com/ntunlp/LLMSanitize.
**footnotetext: Authors contributed equallyfootnotetext: Work done full-time at Salesforce
1 Introduction
In the rapidly evolving landscape of artificial intelligence (AI), Large Language Models (LLMs) have emerged as pivotal tools, driving innovation across a wide spectrum of applications, from natural language processing and automated content creation (Achiam et al., 2023; Ding et al., 2023; Qin et al., 2023; Betker et al., 2023; Ding et al., 2024) to complex decision-making systems and autonomous agents (Wei et al., 2022; Li et al., 2023a; Yang et al., 2023a; Wu et al., 2023). At their core, these models rely on extensive datasets (Raffel et al., 2020; Gao et al., 2020) to learn about language and the world, and generate responses that are increasingly indistinguishable from human-authored output (Ouyang et al., 2022a). However, the integrity of these datasets is paramount, as any contamination—be it intentional manipulation or inadvertent inclusion of biased, inaccurate, or misleading information—can significantly impair the models’ effectiveness and reliability.
Contamination is a broad issue that we divide into data contamination and model contamination. On one hand, data contamination refers to the scenario where an evaluation dataset is (partly) included in the training set, resulting in evaluation results not properly assessing the performance of LLMs. On the other hand, model contamination stands for the phenomenon where the model has seen a specific evaluation dataset during pre-training or fine-tuning, leading to biased results on this dataset. Additionally, if we consider whether the contaminated data contains only inputs or both inputs and labels, contamination includes input contamination and input+label contamination. Depending on the level of access to the model, model contamination detection mainly focuses on white-box detection where we have full access to the model, e.g., a local model, and black-box detection where only the API access is available, e.g., GPT-4 (Achiam et al., 2023).
{forest} for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, text width=2.7cm, s sep=6pt, calign=edge midpoint, , [Contamination, parent [Data Contamination, fill=yellow!30, domain_box [String Matching, fill=orange!30, domain_box [GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), models ] ] [Embeddings Similarity, fill=orange!30, domain_box [Platypus (Lee et al., 2023), Phi-1 (Gunasekar et al., 2023), models ] ] [LLM-based Data Contamination Detection, fill=orange!30, domain_box [LLM Decontaminator (Yang et al., 2023b), Clean-Eval (Zhu et al., 2023), models] ] ] [Model Contamination, fill=yellow!30, domain_box [Performance Analysis, fill=orange!30, domain_box [Task contamination (Li & Flanigan, 2023), Pre-training GPT-2 on contaminated data (Jiang et al., 2024b), models ] ] [Model Completion, fill=orange!30, domain_box [Model Memorization, fill=red!20, domain_box [Injecting test set in training set (Magar & Schwartz, 2022), Completion prompting (Anil et al., 2023), models ] ] [Prompt Engineering, fill=red!20, domain_box [Repetition attack (Nasr et al., 2023), According to prompting (Weller et al., 2023), models ] ] [Model Likelihood, fill=red!20, domain_box [Min-KProb (Shi et al., 2024), Sharded likelihood (Oren et al., 2023), models ] ] ] [LLM-based Model Contamination Detection, fill=orange!30, domain_box [Guided prompting (Golchin & Surdeanu, 2023b), Data Contamination Quiz (Golchin & Surdeanu, 2023a), models] ] ] ]
Figure 1: Classification of the different types of contamination detection methods, with (non-exhaustive lists) examples of methods within each category.
Contamination poses a multifaceted challenge, threatening not only the technical accuracy of LLMs but also their ethical and commercial viability. In contexts where precision and trustworthiness are non-negotiable, such as medical diagnosis, legal advice, or financial services, the repercussions of relying on contaminated data can be profound. Moreover, the allure of leveraging LLM outputs to attract investment underscores a pressing commercial dimension. As businesses increasingly integrate AI-driven insights into their strategic planning and operational decisions, the assurance of data purity becomes intertwined with potential market success and valuation. The credibility of LLM outputs can dramatically influence investor confidence, shaping perceptions of innovation, technological prowess, and, ultimately, financial prospects. Lastly, in the current intense race to build the most powerful LLM (Touvron et al., 2023; Team et al., 2023; Jiang et al., 2024a; Bai et al., 2023; Bi et al., 2024; Young et al., 2024; Gemma Team Google DeepMind, 2024; Chen et al., 2023), the community is struggling to settle on a fixed subset of benchmarks as LLMs performance increases fast (Clark et al., 2018; Zellers et al., 2019; Hendrycks et al., 2020; Lin et al., 2021; Sakaguchi et al., 2021; Cobbe et al., 2021), an issue which is further fueled by underlying contamination.
This landscape necessitates a comprehensive survey and a shared integrated resource of contamination detection in LLMs. We delineate the scope and nature of contamination challenges, identifying their sources, types, and impacts on model performance and applicability. In doing so, we not only contribute to the academic and practical understanding of data contamination issues but also highlight the pressing need for strategies that mitigate these risks, ensuring that the deployment of LLMs aligns with principles of integrity, fairness, and economic sustainability.
To push forward research in contamination detection, we make the following contributions:
  • To the best of our knowledge, we are the first to thoroughly review the field of contamination detection in LLMs.
  • We categorize contamination into two broad types of use cases, namely data and model contamination, and review all existing work in each category. We refer to Figure 1 for an overview of our proposed classification of contamination detection methods.
  • We introduce LLMSanitize, an open-source Python library aimed at helping the community centralize a shared implementation of all data and model contamination methods.
The rest of this paper is structured as follows: Section 2 explores methods and findings related to data contamination. Section 3 addresses the issue of model contamination, where models may be evaluated on data that they have previously seen. Section 4 discusses the current and future challenges in the field, including best practices. Section 5 introduces our library LLMSanitize implementing contamination detection algorithms.
2 Data Contamination
In data contamination detection, given a training dataset $D$ and an evaluation dataset $D_{E}$, the objective is to check whether $D\cap D_{E}$ is empty. In general, checking data contamination is primordial to prevent memorization of contaminated data, which could skew performance on the evaluation benchmarks.
2.1 String Matching
Several LLM training reports assess the level of contamination between their downstream evaluation datasets and their pre-training set through string-matching, the most straightforward data contamination assessment approach. Here we review all such methods.
GPT-2 (Radford et al., 2019) calculates contamination as the percentage of 8-grams from a particular evaluation set that are also found in the WebText training set. As a result, they discover 1-6% overlap between common LM datasets’ test sets and the WebText dataset used for pre-training, with an average of 3.2%. This method also exposes contamination between test sets and training sets of these datasets.
GPT-3 (Brown et al., 2020) defines a clean version of each evaluation dataset by removing all data points having a 13-gram collision with anything in the pre-training Common Crawl (C4) dataset (Raffel et al., 2020). A conservative overlap analysis discovers large contamination problems for common datasets, showing some Wikipedia language modeling benchmarks, the Children’s Book Test dataset (Hill et al., 2015), Quac (Choi et al., 2018) and SQuAD 2.0 (Rajpurkar et al., 2018) to be nearly entirely contaminated.
Dodge et al. (2021) investigate the C4 (Raffel et al., 2020) dataset and study benchmark data contamination, which measures how much training or test datasets from downstream NLP tasks appear in the pre-training corpus. After studying both input-label and input contamination, they discover varied contamination, ranging from less than 2% to over 50%.
PaLM (Chowdhery et al., 2023) identifies 10 evaluation datasets at risk of contamination per their construction process, and for each dataset, partitions it into a clean and a contaminated subset based on whether at least 70% of the 8-grams of the data point can be found at least once in the training set. Like GPT-3, they conclude that the performance gap between clean sets and contaminated sets is mostly small across sets of various contamination degrees.
Since then, GPT-4 (Achiam et al., 2023) measures cross-contamination between evaluation sets and pre-training data with substring match: a match is identified if any of three 50-characters randomly sampled substrings from the evaluation data point (or the entire data point if it smaller than this) is a substring of the training set. After examining GPT-4 performance gap between contaminated and clean subsets, the authors conclude that contamination overall has very little effect on the reported zero-shot results.
Llama-2 (Touvron et al., 2023) refines the methods above by using tokens instead of words to measure contamination, and establishes that a token is contaminated if it appears in any token n-gram longer than 10 tokens in both the evaluation sample and the training set. From there, the contamination level of a sample is defined as its percentage of contaminated tokens, yielding a granular definition of data contamination.
Li (2023b) computes the METEOR (Banerjee & Lavie, 2005) score between matched pages from CommonCrawl and queries from Bing API and considers those with scores over 0.75 as contaminated. Varying contamination levels are identified, ranging from 1% to 47% across popular benchmarks. Similar to PaLM and GPT-4, the study also finds that data contamination does not necessarily lead to increased performance metrics.
Deng et al. (2023) propose to detect contamination by retrieving top-10 documents from pre-training datasets The Pile (Gao et al., 2020) or C4 (Raffel et al., 2020), then split retrieved documents into 13-gram chunks, and compute overlap metrics between these chunks and the evaluation data point’s chunks. Results show that TruthfulQA exhibits high overlap with the pre-training datasets.
2.2 Embeddings Similarity
Beyond simple string matching, which fails on paraphrased samples, computing cosine similarity between embeddings offers an attractive alternative, as this process is more robust to vocabulary changes, including paraphrases or synonyms.
Lee et al. (2023) prevent contamination in their newly introduced Open-Platypus dataset by removing test questions which have a cosine similarity (using SentenceTransformer embeddings (Reimers & Gurevych, 2019)) greater than 80% against any training item.
Phi-1 (Gunasekar et al., 2023) runs a data contamination analysis between their fine-tuning set CodeExercises and the evaluation set HumanEval (Chen et al., 2021). They show that embeddings-based retrieval between code snippets using L2 distance between normalized CodeGen-Mono 350M embeddings is effective, whereas n-gram overlap fails in the coding domain as it cannot capture the similarities in the logic between two coding functions.
The work from (Riddell et al., 2024) provides a hybrid approach combining string matching and embeddings similarity: an aggregated score is designed as the maximum between the Levenshtein edit distance and the similarity from the Dolos toolkit (Maertens et al., 2022). They show that widely used code generation benchmarks MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021) are contaminated with pre-training sets The Pile (Gao et al., 2020) and especially The Stack (Kocetkov et al., 2022), which significantly inflates performance of models like CodeGen-NL (Nijkamp et al., 2022).
2.3 LLM-based data contamination detection
String-matching could struggle to detect synthetic training data produced by LLMs. It is also challenging to choose an appropriate similarity threshold in semantic similarity search approaches using embeddings. To improve robustness, some methods attempt to use LLMs themselves to detect contamination, leveraging their high performance at paraphrasing and paraphrase detection (Witteveen & Andrews, 2019; Abaskohi et al., 2023).
Yang et al. (2023b) argue that contamination checks become challenging because of the presence of “rephrased samples”, which have the same semantics but different surface form as the original sample. Therefore, they first use embeddings similarity search to get the top-k similar samples with a given test sample, then prompt a strong LLM (namely, GPT-4) to examine whether any of the top-k samples is too close to the test case. Results show that this LLM Decontaminator algorithm works significantly better than existing methods at tagging contaminated samples.
Zhu et al. (2023) introduce Clean-Eval, a benchmark to assess the decontaminated performance of LLMs. Data points from popular benchmarks (e.g., MMLU) are paraphrased with GPT-3.5, and the LLM is evaluated on the paraphrased samples with the lowest BLEURT score with regards to the original ground truth. Results show a drop in performance for GPT-3.5 and Llama-2 on the Clean-Eval data, confirming prior contamination.
3 Model Contamination
In model contamination, a model $M$ trained on a training set $D_{M}$ is applied to an evaluation dataset $D_{E}$ and the task is to determine whether $D_{M}\cap D_{E}$ is empty. This exposure to evaluation data during training raises concerns about the model’s ability to generalize to unseen data, as its performance could be artificially inflated. Compared to data contamination, the training set is dependent on the model, and could be unknown, as with GPT-3.5 and GPT-4.
3.1 Performance Analysis
We first review a line of research assessing model contamination through performance analysis: $M$ is applied to evaluation datasets of different timestamps, including recent datasets which did not exist at pre-training time, forfeiting any contamination. A drop in performance on these recent data points suggests previous contamination.
Roberts et al. (2023) introduce a comparative analysis of the efficacy in code generation tasks between GPT-4 and GPT-3.5-Turbo, utilizing datasets from periods before and after a specified cut-off date. Through a meticulous examination of empirical evidence, the study reveals a compelling correlation between the likelihood of exposure to coding problem platforms such as Codeforces and Project Euler by these LLMs and the observed test case pass rates. Specifically, the analysis underscores a statistically significant positive correlation for problems catalogued on GitHub prior to the GPT models’ training cut-off, suggesting that their inclusion in the training data enhanced their performance on related tasks.
Li & Flanigan (2023) present four innovative methods to detect task contamination in LLMs which are crucial for maintaining the integrity of machine learning processes. These methods include (i) training data inspection to identify direct task examples within the data, (ii) task example extraction from instruction-tuned models, (iii) membership inference to detect exact content matches suggesting data inclusion, and (iv) chronological analysis to assess contamination through performance metrics against known dataset timelines. This multifaceted approach provides a comprehensive toolkit for researchers and developers to safeguard against inadvertent task leakage, thereby enhancing LLMs’ reliability.
Jiang et al. (2024b) conduct a comprehensive evaluation of data contamination in language model pre-training through experiments on various public datasets. The research highlights the critical issue of evaluation data leakage and its influence on language model performance, utilizing a novel approach by pre-training models from scratch. Findings reveal a U-shaped relationship between the extent of data contamination and model performance, challenging existing n-gram based definitions of contamination as inadequate. This underscores the necessity for more robust contamination detection methods.
3.2 Model Completion
An intuitive approach to detect model contamination in LLMs is to carefully analyze their outputs in controlled prompting setups. Outputs on contaminated data points may be too close or exactly match training data, or be subject to very high model confidence.
3.2.1 Model Memorization
When analyzing model memorization, researchers focus on which training data points have been entirely memorized by the model, a process that can readily be leveraged to flag contamination on downstream tasks.
Elangovan et al. (2021) highlight that the common practice of random data shuffling in NLP tasks fails to account for the overlaps between training and test sets, leading to data leakage and inflated performance metrics. Models that merely memorize instead of generalize may benefit from this overlap. By employing a bag-of-words approach to assess text similarity, the study provides a foundational method for identifying and mitigating leakage, though it acknowledges the complexity of semantic similarity measurement and the potential for more sophisticated methodologies to refine overlap detection.
Lee et al. (2021) discuss the substantial benefits of deduplicating datasets for training. Deduplication significantly reduces the incidence of memorized data emission by a factor of 10×, mitigates the common issue of train-test overlap, thereby preventing overestimation of model accuracy, enhances training efficiency by reducing dataset size by up to 19%, and maintains or improves model perplexity, allowing for faster attainment of higher accuracy. This comprehensive approach not only streamlines the training process but also contributes to more environmentally and economically sustainable model development, without compromising data quality or model performance.
Magar & Schwartz (2022) find evidence of contamination across several benchmarks, impacting the performance of models like GPT-3. A novel approach to quantify the effect of contamination involves training a model on a mix of general and task-specific data, then comparing its performance on seen versus unseen instances to measure memorization and exploitation. This methodology highlights the nuanced relationship between memorization and exploitation, indicating that larger models or those trained with certain data configurations may exhibit different levels of sensitivity to contaminated data.
Gemma Team Google DeepMind (2024) emphasize the potential vulnerabilities of aligned models to adversarial attacks inducing recitation of memorized training data. Through comparable methods to Anil et al. (2023), they analyze discoverable memorization across 10k documents from the Gemma pre-trained models’ corpora. The analysis distinguishes between exact memorization—where the model’s generated text precisely matches the source—and approximates memorization, gauged by a 10% edit distance threshold. The findings reveal that Gemma models exhibit memorization rates comparable to those of the PaLM and PaLM-2 models.
3.2.2 Prompting Techniques
Researchers have engineered elaborated prompting techniques that elicit data point completion by the LLM, where an output suspiciously close to the actual training data point could indicate contamination. Standard prompting asking the LLM to complete a data point may fail due to the LLM’s alignment during RLHF (Ouyang et al., 2022b).
Nasr et al. (2023) uncover the vulnerability of LLMs including ChatGPT to data extraction attacks, revealing a critical privacy concern. By prompting with an off-(training)-distribution input, they show that ChatGPT can regurgitate entire chunks of training data, including sensitive personal-level information. The study illustrates that even models trained with alignment techniques, aimed at reducing the emission of memorized data, can be induced to divulge substantial amounts of sensitive information, calling for the development of more robust defenses. The implications of this research are far-reaching, emphasizing the importance of addressing the privacy vulnerabilities inherent in the deployment of LLMs.
Weller et al. (2023) highlight the potential for steering LLMs towards generating more factual and grounded content through effective prompting strategies. By appending instructional phrases such as Accoding to that encourage quoting from specific corpora, this study observes improvements in grounding as measured by the Quoted Information Precision (QUIP) Score. This method shows promise across various domains and corpora, indicating its versatility and effectiveness in leveraging LLMs’ memorized knowledge for generating more accurate and reliable responses.
3.2.3 Model Likelihood
Methods based on the likelihood utilize the fact that a model’s next token predictions are more likely to be more confident if the data point has been seen during training.
The Min-KProb technique (Shi et al., 2024) assumes white-box access to the LLM’s logits or next-token probability distributions. Given a test sample $X$, the method consists in running the LLM through all tokens of $X$, and keeping track of the $k$tokens with the smallest predicted probability. Then, the average between such bottom probabilities is computed and $X$ is deemed contaminated if this average is too high.
Oren et al. (2023) take an original approach to analyze the LLM’s likelihood on an evaluation dataset $D_{E}$. They run inference on $D_{E}$ and on shuffled versions of $D_{E}$. LLM log probabilities being statistically different on the non-shuffled dataset compared to the shuffled versions signifies contamination. The method relies on the assumption that if they are included in the pre-training set, evaluation datasets tend to be present with the same default ordering.
Li (2023a) compares perplexity on benchmark samples against memorized and clean baselines. The study finds significant memorization in recent models on popular reading comprehension and summarization benchmarks, while multiple-choice benchmarks show less evidence of contamination. This method provides a tool for the community to conduct rigorous contamination analysis, enabling more accurate and reliable model evaluation.
Dong et al. (2024) propose two novel likelihood-based contamination detection methodologies: CDD (Contamination Detection via output Distribution) and TED (Trustworthy Evaluation via output Distribution). CDD detects data contamination by observing the peakedness in the LLM’s output distribution in a black-box manner. It represents a significant improvement over existing approaches, offering average relative improvements of 21.8%-30.2% in terms of Accuracy, F1 Score, and AUC metrics. TED corrects the LLM’s output distribution to mitigate the impact of data contamination on evaluation metrics, significantly reducing performance improvements attributed to data contamination across various scenarios and contamination degrees.
3.3 LLM-based Model Contamination Detection
LLMs also establish a new paradigm in model contamination detection.
Golchin & Surdeanu (2023b) introduce guided prompting: unlike a standard prompt asking the LLM to complete the data point, guided prompting also includes extra information such as the dataset name. Contamination is then assessed based on the averaged difference in performance between standard and guided prompting, or if GPT-4 with in-context learning finds an exact match or two near-exact matches in the guided completions. This latter method with GPT-4 is highly accurate in identifying contamination (92%–100%). Moreover, the investigation highlights the prevalence of contamination in datasets such AG News (Zhang et al., 2015), WNLI (Levesque et al., 2012), and XSum (Narayan et al., 2018), emphasizing the critical need to address data integrity in LLM applications.
The same authors follow up with the Data Contamination Quiz (DCQ) evaluation framework for contamination detection in black-box LLMs (Golchin & Surdeanu, 2023a). The LLM under investigation is prompted with five completion options, where one option is the ground-truth text from the original dataset, three other options are paraphrases from GPT-4, and the last choice is None. The assumption is that if the LLM picks the exact answer, it is doing so out of memorization, and the authors show that this DCQ framework finds contamination cases than the guided prompting method.
4 Discussion
4.1 Best Practices
Recent works have called the community to adopt better practices to reduce contamination.
Encrypting evaluation datasets
Jacovi et al. (2023) advocate for the encryption of publicly released test data using a public key, coupled with licensing agreements that prohibit derivative distribution. They recommend implementing training exclusion controls for closed APIs and safeguarding test data by refusing to perform evaluations without such controls. Moreover, they suggest avoiding data that appears with the solution and releasing internet-derived data along with the corresponding web-page context.
Scanning newly released evaluation datasets
Sainz et al. (2023) argue to develop automatic or semi-automatic approaches to detect contamination for new benchmarks and design mechanisms to flag those works with contamination.
Evaluating on a wide spectrum
To comprehensively and fairly evaluate LLMs, Zhou et al. (2023) suggest to use more benchmarks from diverse sources, covering both basic and advanced capabilities. Furthermore, they recommend employing multiple task prompts for benchmark tests to derive model performance that is more stable and reliable. Additionally, they highlight the necessity of conducting data decontamination checks between pre-training data and any evaluation data when using benchmarks.
Not leaking data to closed-source APIs
Aside from technical methodologies, Balloccu et al. (2024) conduct a systematic literature review of 225 papers and carefully detail data leakage from them to closed-source models like the GPT-series. Overall, they conclude that $\sim$42of reviewed papers leaked data to GPT-3.5 and GPT-4 for a total of $\sim$4.7M benchmark samples across 263 benchmarks. The authors also report a series of evaluation malpractices and propose a list of suggested practices for evaluating closed-source LLMs.
4.2 New Evaluation Benchmarks
Several datasets have been introduced aiming to provide contamination-free evaluation:
  • LatestEval (Li et al., 2023b) leverages the most recent texts to create dynamic reading comprehension evaluations. The benchmark is constructed in three steps: (i) collect the latest texts; (ii) extract key information, and (iii) construct questions based on the extracted information through template-filling or LLMs.
  • WIKIMIA (Shi et al., 2024) is a dynamic benchmark made of Wikipedia events created after model training (i.e., after 2023). They further introduce the paraphrase setting where ChatGPT is leveraged to create paraphrased examples for evaluation, as also performed in CleanEval (Zhu et al., 2023).
  • KIEval (Yu et al., 2024) is an interactive evaluation framework incorporating an LLM-powered ”interactor”, which can ask follow-up questions in multi-round dialogue that leads to contamination-resilient evaluation. The evaluator could eventually determine whether a model’s response is merely a recall of benchmark answers or demonstrates a deep comprehension to apply knowledge in more complex conversations.
  • LiveCodeBench (Jain et al., 2024) continuously collects new problems from LeetCode, AtCoder and CodeForces to provide a contamination-free code generation benchmark. They show a drop in performance on new LeetCode problems for some models including DeepSeek (Bi et al., 2024), hinting at contamination.
  • Termite (Ranaldi et al., 2024) is a new text-to-SQL dataset, locked out of public access through search engines via an encryption key. Termite is made of handcrafted databases each paired with around five queries, designed to match properties of the Spider dataset (Yu et al., 2018), which unfortunately shows high signs of contamination on GPT-3.5.
Moreover, Alzahrani et al. (2024) point out that slight perturbations to existing benchmarks may disrupt model rankings on leaderboards, calling for better evaluation benchmarks.
LLM Average ARC HellaSwag MMLU TruthulQA Winogrande GSM8K
Llama-2-7B 0.278 0.064 0.489 0.146 0.001 0.001 0.968
Qwen-1.5-7B 0.432 0.000 0.023 0.700 0.664 0.202 1.000
Mistral-7B 0.312 0.000 0.000 0.010 0.042 0.999 0.444
Gemma-7B 0.417 0.978 0.000 0.580 0.122 0.820 0.000
[…]
Table 1: P-value of the guided prompting (with ROUGE-L) contamination detection (Golchin & Surdeanu, 2023b) for five LLMs on OpenLLM Leaderboard datasets. Shaded results are not statistically significant.
4.3 Future Challenges
Given the fast and ever-changing landscape of machine learning research, future directions for contamination detection in LLMs could encompass a broad variety of methodologies and technological aspects. We highlight a few potentially critical areas of focus:
Real-time contamination detection systems
Real-time data contamination detection systems that can monitor data streams continuously and alert users to potential contamination events as they happen would be of upmost importance in critical applications like finance (Frésard et al., 2011) and healthcare (Gao et al., 2022) where data integrity is paramount and model reliability critical. Given the sheer volume of new data uploaded daily on the Internet, solving this task requires major technological breakthroughs.
Bypassing contamination detection
Dekoninck et al. (2024) show a very effective way to bypass several existing contamination detection methods: Evasive Augmentation Learning (EAL) consists in paraphrasing benchmarks with GPT-4 and fine-tuning the LLM on the paraphrased data. This method can bypass even oracle-access contamination detection methods, calling for the development of more robust methods.
Ethical and legal data frameworks
We believe in the need for comprehensive ethical and legal frameworks (Chang, 2021) to govern the collection, usage, and management of data for LLM training. This includes policies and protocols for data privacy, consent, and use that help prevent the incorporation of contaminated data from unethical sources and the contamination of widely used pre-training data sources (e.g, CommonCrawl). Developing contamination detection techniques that do not compromise individual privacy will be crucial (Lin et al., 2010; Hayes & Ohrimenko, 2018).
5 The LLMSanitize Library
LLM Average ARC HellaSwag MMLU TruthulQA Winogrande GSM8K BBC 2024
Llama-2-7B 0.389 0.258 0.000 0.967 0.154 0.171 0.781 0.889
Qwen-1.5-7B 0.247 0.426 0.000 0.476 0.255 0.115 0.209 0.698
Mistral-7B 0.381 0.862 0.000 0.376 0.057 0.485 0.507 0.304
Gemma-7B 0.300 0.234 0.000 0.255 0.450 0.215 0.648 0.371
Table 2: P-value of the sharded likelihood test (Oren et al., 2023) for four LLMs on OpenLLM Leaderboard datasets. Shaded results are not statistically significant.
LLM Average ARC HellaSwag MMLU TruthulQA Winogrande GSM8K BBC 2024
Llama-2-7B 9.01 8.92 8.84 9.06 9.27 10.88 7.07 6.98
Qwen-1.5-7B 10.58 10.69 10.10 10.47 11.20 14.02 7.02 8.39
Mistral-7B 9.76 9.77 10.49 9.53 9.26 12.60 6.91 6.34
Gemma-7B 76.23 53.43 33.47 59.27 157.82 125.84 27.55 18.48
Table 3: Min-20% Prob (Shi et al., 2024) for four LLMs on OpenLLM Leaderboard datasets. Lower values mean greater risk of contamination.
To facilitate progress in contamination detection in LLMs, we have built and are releasing an open-source Python library, named LLMSanitize. LLMSanitize supports both data contamination and model contamination types of use cases, and input/label contamination. A minimal pseudo-code to use our library is of the following structure:
from llmsanitize import DataContaminationChecker, ModelContaminationChecker args = <…> contamination_checker = ModelContaminationChecker(args) contamination_checker.run_contamination(args.method)
The user is expected to specify the training and evaluation datasets for data contamination ; and the LLM and evaluation dataset for model contamination. Datasets and LLMs are expected to be available on HuggingFace hub, namely transformers (Wolf et al., 2020) and datasets (Lhoest et al., 2021) libraries, which our code builds upon. Efficient LLM inference for model-based use cases is handled with the vLLM library (Kwon et al., 2023). Other key parameters being controlled through arguments include the number of evaluation data points to process, and hyper-parameters specific to each contamination detection method.
We now put to test LLMSanitize. We apply three recent popular model contamination methods which we reviewed above: guided prompting (Golchin & Surdeanu, 2023b), sharded likelihood (Oren et al., 2023) and Min-KProb (Shi et al., 2024) on four widely used 7B-size LLMs: Llama-2 (Touvron et al., 2023), Qwen-1.5 (Bai et al., 2023), Mistral (Jiang et al., 2023) and Gemma (Gemma Team Google DeepMind, 2024). We use the chatbot version for each LLM, and measure contamination on the six datasets of HuggingFace OpenLLM Leaderboard: ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2020), TruthulQA (Lin et al., 2021), Winogrande (Sakaguchi et al., 2021) and GSM8K (Cobbe et al., 2021) ; subsampling 100 data points from each test set. We also evaluate GPT-3.5-turbo on guided prompting - the blocked access to the logprobs prohibits us from applying the other two methods. Lastly, for comparison purpose we report likelihood-based results on contamination-free data: BBC news articles from March 2024.
Results for guided prompting, sharded likelihood and Min-KProb are in Table 1, Table 2, and Table 3, respectively. The HellaSwag benchmark (Zellers et al., 2019) shows strong signs of contamination, consistent across methods and LLMs. We also report data contamination results in Appendix A and model contamination results with base models in Appendix B.
6 Conclusion
As LLMs rapidly evolve and the data used for training become more extensive, LLMs’ performances are inevitably biased by contamination. Therefore, in this survey, we investigated and structured the current landscape of contamination detection in LLMs. We introduced a new Python library named LLMSanitize which is a universal tool for quickly detecting contamination in various models and datasets. The extensive investigation and proposed tool could provide a universal foundation for future models when considering data or model contamination issues.
Appendix A Data Contamination Detection Results
Data contamination detection method ARC HellaSwag Winogrande GSM8K
Valid Test Valid Test Valid Test Test
GPT-2 style 10.70 11.26 13.88 13.97 0.24 0.62 11.30
GPT-3 style 0.67 0.51 0.62 0.92 0.00 0.11 2.73
Exact match 0.00 0.00 0.00 0.00 0.00 0.00 0.00
PaLM-style 0.33 0.43 0.00 0.00 0.00 0.00 0.00
GPT-4 style 0.67 0.43 0.19 0.38 0.00 0.11 0.30
Platypus-style 11.37 9.56 22.14 20.08 0.87 1.41 7.51
Table 4: Data contamination detection results for several baseline methods described in Section 2. We compare each dataset validation or test set against its training set, and report the percentage of data points flagged as contaminated. All results were obtained using the LLMSanitize library, without subsampling any dataset partition.
In Table 4, we benchmark several data contamination techniques on datasets from the OpenLLM Leaderboard following Tables 1, 3 and 2. We discard datasets which do not have a training set (MMLU and TruthfulQA), and perform data contamination analysis between the training set and the validation (if available) and test sets. We notice that ARC and HellaSwag show on average higher contamination with the training set, especially with GPT-2-style string matching and Platypus-style embeddings similarity.
Appendix B More Model Contamination Detection Results
LLM Average ARC HellaSwag MMLU TruthulQA Winogrande GSM8K
Llama-2-7B-base 0.566 0.001 0.847 0.690 0.913 0.945 0.000
Llama-2-7B 0.278 0.064 0.489 0.146 0.001 0.001 0.968
[…]
Qwen-1.5-7B 0.432 0.000 0.023 0.700 0.664 0.202 1.000
[…]
Mistral-7B 0.312 0.000 0.000 0.010 0.042 0.999 0.444
[…]
Gemma-7B 0.417 0.978 0.000 0.580 0.122 0.820 0.000
Table 5: P-value of the guided prompting (with ROUGE-L) contamination detection (Golchin & Surdeanu, 2023b) for four open-source LLMs in both their base and instruction-tuned versions on OpenLLM Leaderboard datasets. Shaded results are not statistically significant.
In Table 5, we reproduce the guided prompting contamination detection results from Table 1 while including the base version (immediately after pre-training) of each LLM. Some datasets such as ARC or TruthfulQA seem to have contaminated all pre-training datasets. We note that although they are consistent, contamination levels may change in both directions from pre-training to instruction-tuning: for instance Llama-2 becomes contaminated on TruthfulQA and Winogrande after instruction tuning ; while Gemma does not show contamination anymore after instruction tuning on MMLU or TruthffulQA. We highlight that designing contamination detection techniques robust to the changes in the LLM knowledge after undergoing instruction tuning and alignment remains an open research problem.