### v3

# How Much are Large Language Models Contaminated? A Comprehensive Survey and the LLMSanitize Library

Abstract
With the rise of Large Language Models (LLMs) in recent years, abundant new opportunities are emerging, but also new challenges, among which contamination is quickly becoming critical. Business applications and fundraising in AI have reached a scale at which a few percentage points gained on popular question-answering benchmarks could translate into dozens of millions of dollars, placing high pressure on model integrity. At the same time, it is becoming harder and harder to keep track of the data that LLMs have seen; if not impossible with closed-source models like GPT-4 and Claude-3 not divulging any information on the training set. As a result, contamination becomes a major issue: LLMs’ performance may not be reliable anymore, as the high performance may be at least partly due to their previous exposure to the data. This limitation jeopardizes the entire progress in the field of NLP, yet, there remains a lack of methods on how to efficiently detect contamination. In this paper, we survey all recent work on contamination detection with LLMs, and help the community track contamination levels of LLMs by releasing an open-source Python library named LLMSanitize implementing major contamination detection algorithms, available here: https://github.com/ntunlp/LLMSanitize
How Much are Large Language Models Contaminated? A Comprehensive Survey and the LLMSanitize Library
Mathieu Ravaut1,2, Bosheng Ding1, Fangkai Jiao1,2, Hailin Chen1,3, Xingxuan Li1, Ruochen Zhao1, Chengwei Qin1, Caiming Xiong3, Shafiq Joty1,3† 1 Nanyang Technological University, Singapore 2 Institute of Infocomm Research (I2R), ASTAR, Singapore 3 Salesforce Research {mathieuj001,bosheng001,fangkai002,hailin001,xingxuan001,ruochen002,chengwei003}@e.ntu.edu.sg {cxiong,sjoty}@salesforce.com
**footnotetext: Authors contributed equallyfootnotetext: Work done full-time at Salesforce
1 Introduction
In the rapidly evolving landscape of artificial intelligence (AI), Large Language Models (LLMs) have emerged as pivotal tools, driving innovation across a wide spectrum of applications, from natural language processing (NLP) and automated content creation (Achiam et al., 2023; Betker et al., 2023) to complex decision-making systems and autonomous agents (Wei et al., 2022; Li et al., 2023a; Yang et al., 2023a; Wu et al., 2023). At their core, these models rely on extensive datasets (Raffel et al., 2020; Gao et al., 2020) to learn about language and the world, and generate responses that are increasingly indistinguishable from human-authored output (Ouyang et al., 2022a). However, the integrity of these datasets is paramount, as any contamination can significantly impair the models’ effectiveness and reliability.
Detecting contamination in a LLM is a broad issue that we divide into open-data contamination detection and closed-data contamination detection. On the one hand, open-data refers to the scenario where the LLM pre-training data is known, enabling direct comparisons with the evaluation dataset. On the other hand, closed-data contamination refers to the more and more prevalent use case of an unknown pre-training set. This latter case is more challenging and necessitates an examination of the LLM’s behavior on evaluation data points. Additionally, we highlight the following orthogonal dimension to the duality explained above: if we consider whether the contaminated data contains only inputs or both inputs and labels, contamination includes input contamination and input+label contamination. A third dimension to the problem is the level of access to the LLM itself in closed-data setups: white-box detection where we have full access to the model, e.g., a local model, and black-box detection where only the API access is available, e.g., GPT-4 (Achiam et al., 2023).
{forest} for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, text width=2.7cm, s sep=6pt, calign=edge midpoint, , [Contamination
Detection Methods, parent [Open-Data
Contamination Detection, fill=yellow!30, domain_box [String
Matching, fill=orange!30, domain_box [GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), PaLM (Anil et al., 2023), GPT-4 (Achiam et al., 2023), Llama-2 (Touvron et al., 2023), Dodge et al. (2021); Li (2023b); Deng et al. (2023), models ] ] [Embeddings
Similarity, fill=orange!30, domain_box [Platypus (Lee et al., 2023), Phi-1 (Gunasekar et al., 2023), Riddell et al. (2024), models ] ] [Model-based Open-
Data Contamination Detection, fill=orange!30, domain_box [LLM Decontaminator (Yang et al., 2023b), models] ] ] [Closed-Data
Contamination Detection, fill=yellow!30, domain_box [Performance
Analysis, fill=orange!30, domain_box [Roberts et al. (2023); Li and Flanigan (2023); Jiang et al. (2024b), models ] ] [Model
Memorization, fill=orange!30, domain_box [TS-Guessing (Deng et al., 2023), Elangovan et al. (2021); Magar and Schwartz (2022); Gemma Team Google DeepMind (2024); Nasr et al. (2023); Weller et al. (2023); Chang et al. (2023); Ranaldi et al. (2024), models ] ] [Model
Confidence, fill=orange!30, domain_box [Min-KProb (Shi et al., 2024), CDD and TED (Dong et al., 2024), Li (2023a); Oren et al. (2023); Xu et al. (2024b), models ] ] [Model-based Closed-
Data Contamination Detection, fill=orange!30, domain_box [Guided prompting (Golchin and Surdeanu, 2023b), Tu et al. (2024), models] ] ] ]
Figure 1: Classification of the different types of contamination detection methods.
Contamination poses a multifaceted challenge, threatening not only the technical accuracy of LLMs but also their ethical and commercial viability. In high-stakes scenarios where imprecision can have dramatic consequences, such as medical diagnosis, legal advice, or financial services, the repercussions of relying on contaminated data can be profound. Moreover, the allure of leveraging LLM outputs to attract investment underscores a pressing commercial dimension. As businesses increasingly integrate AI-driven insights into their strategic planning and operational decisions, the assurance of data purity becomes intertwined with potential market success and valuation. Lastly, in the current intense race to build the most powerful LLM (Touvron et al., 2023; Team et al., 2023; Jiang et al., 2024a; Bai et al., 2023; Bi et al., 2024; Young et al., 2024; Gemma Team Google DeepMind, 2024; Chen et al., 2023), the community is struggling to settle on a fixed subset of benchmarks as LLMs performance increases fast (Clark et al., 2018; Zellers et al., 2019; Hendrycks et al., 2020; Lin et al., 2021; Sakaguchi et al., 2021; Cobbe et al., 2021), an issue which is further fueled by underlying contamination.
This landscape necessitates a comprehensive survey and a shared integrated resource of contamination detection in LLMs. We delineate the scope and nature of contamination detection challenges, identifying their types. In doing so, we not only contribute to the academic and practical understanding of data contamination issues but also highlight the pressing need for strategies that mitigate these risks.
To push forward research in contamination detection, we make the following contributions:
  • To the best of our knowledge, we are the first to thoroughly review the field of contamination detection in LLMs.
  • We categorize contamination detection into two broad types of use cases vastly differing in the techniques involved, namely open-data and closed-data contamination detection, and review all existing work in each category. We refer to Figure 1 for an overview of our proposed classification of contamination detection methods.
  • We introduce LLMSanitize, an open-source Python library aimed at helping the community centralize a shared implementation of all contamination detection methods.
The rest of this paper is structured as follows: Section 2 explores methods and findings related to open-data contamination detection. Section 3 addresses the issue of closed-data contamination detection, where models may be evaluated on data that they have previously seen. Section 4 discusses the current and future challenges in the field, including best practices. Section 5 briefly introduces our library LLMSanitize implementing contamination detection algorithms, which is further described in Appendix.
2 Open-Data Contamination Detection
In the open-data setup, we assume that the (pre)-training set is known. Given a training dataset $D$ and an evaluation dataset $D_{E}$, the objective is to check whether $D\cap D_{E}$ is empty. In the following, we review existing open-data contamination detection methods, which directly compare ensembles of data points.
2.1 String Matching
The most straightforward method to compute overlap between textual datasets is string matching or analyzing the set of words intersecting two texts.
Several LLM pre-training reports assess contamination levels by using string matching techniques. GPT-2 (Radford et al., 2019) calculates contamination as the percentage of 8-grams from a particular evaluation set that are also found in the WebText training set. An average of 3.2% overlap between common LM datasets’ test sets and the WebText dataset used for pre-training is discovered. GPT-3 (Brown et al., 2020) scans data points having a 13-gram collision with anything in the pre-training Common Crawl (C4) dataset (Raffel et al., 2020). Some Wikipedia language modeling benchmarks, the Children’s Book Test dataset (Hill et al., 2015), Quac (Choi et al., 2018) and SQuAD 2.0 (Rajpurkar et al., 2018) are nearly entirely contaminated. PaLM (Chowdhery et al., 2023) identifies 10 evaluation datasets at risk of contamination per their construction process, and for each dataset, partitions it into a clean and a contaminated subset based on whether at least 70% of the 8-grams of the data point can be found at least once in the training set. Both GPT-3 and PaLM authors conclude that the performance gap between clean sets and contaminated sets is mostly small across sets of various contamination degrees. GPT-4 (Achiam et al., 2023) measures cross-contamination between evaluation sets and pre-training data by computing 50-characters substring collisions. After examining GPT-4 performance gap between contaminated and clean subsets, the authors conclude that contamination overall has very little effect on the reported zero-shot results. Llama-2 (Touvron et al., 2023) refines the methods above by using tokens instead of words to measure contamination, and asserts that a token is contaminated if it appears in any token n-gram longer than 10 tokens in both the evaluation sample and the training set. From there, the contamination level of a sample is defined as its percentage of contaminated tokens.
Other studies analyze overlap between training sets and evaluation sets to assess contamination levels. Dodge et al. (2021) investigate the C4 (Raffel et al., 2020) dataset and study benchmark data contamination, which measures how much training or test datasets from downstream NLP tasks appear in the C4 pre-training corpus. After studying both input-label and input contamination, they discover varied contamination, ranging from less than 2% to over 50%. Li (2023b) compute the METEOR (Banerjee and Lavie, 2005) score between matched pages from CommonCrawl and queries from Bing API and consider those with scores over 0.75 as contaminated. Contamination levels ranging from 1% to 47% across popular benchmarks are identified. Deng et al. (2023) propose to detect contamination by retrieving top-10 documents from pre-training datasets The Pile (Gao et al., 2020) or C4 (Raffel et al., 2020), then splitting retrieved documents into 13-gram chunks, and computing overlap metrics between these chunks and the evaluation data point’s chunks. TruthfulQA (Lin et al., 2021) exhibits high overlap with the pre-training datasets.
2.2 Embeddings Similarity
Beyond simple string matching, computing cosine similarity between embeddings offers an attractive alternative, as this process is more robust to surface-form vocabulary changes like paraphrases.
Lee et al. (2023) prevent contamination in their Open-Platypus dataset by removing test questions which have a cosine similarity (using SentenceTransformer embeddings (Reimers and Gurevych, 2019)) greater than 80% against any training item. Phi-1 (Gunasekar et al., 2023) runs a data contamination analysis between their CodeExercises dataset and the evaluation set HumanEval (Chen et al., 2021). They show that embeddings-based retrieval between code snippets using L2 distance between normalized CodeGen-Mono 350M embeddings is effective, whereas n-gram overlap fails in the coding domain as it cannot capture the similarities in the logic between two coding functions.
The work from (Riddell et al., 2024) provides a hybrid approach combining string matching and embeddings similarity: an aggregated score is designed as the maximum between the Levenshtein edit distance and the similarity from the Dolos toolkit (Maertens et al., 2022). Widely used code generation benchmarks MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021) are contaminated with pre-training sets The Pile (Gao et al., 2020) and especially The Stack (Kocetkov et al., 2022), which significantly inflates performance of models like CodeGen-NL (Nijkamp et al., 2022).
2.3 Model-based Open-Data Contamination Detection
Embeddings similarity is limited by its need to choose an appropriate similarity threshold. To improve robustness, some methods attempt to use LLMs themselves to detect contamination in LLMs, leveraging their high performance in paraphrasing and paraphrase detection (Witteveen and Andrews, 2019; Abaskohi et al., 2023).
Yang et al. (2023b) argue that contamination checks become challenging because of the presence of “rephrased samples”, which have the same semantics but different surface form as the original sample. Therefore, they first use embeddings similarity search to get the top-k similar samples with a given test sample, then prompt a strong LLM (namely, GPT-4) to examine whether any of the top-k samples is too close to the test case. Results show that this LLM Decontaminator algorithm works significantly better than existing methods at tagging contaminated samples.
3 Closed-Data Contamination Detection
In closed-data contamination detection, the training data is unknown. Formally, a model $M$ has been trained on a training set $D_{M}$ (the pre-training set is specific to each LLM) and is applied to an evaluation dataset $D_{E}$. $D_{M}$ is not accessible, and researchers have to use tools analyzing $M$’s behavior on $D_{E}$.
3.1 Performance Analysis
We first review a line of research assessing model contamination through simple performance analysis. $M$ is applied to different evaluation datasets, and a sudden drop in performance suggests previous contamination.
The most straightforward performance analysis consists in plotting performance against time, which leverages the fact that evaluation datasets released after a LLM is pre-trained should not be contaminated. In the code domain, Roberts et al. (2023) investigate the performance of GPT-3.5 and GPT-4 on datasets released before and after the training cutoff for both models. The analysis underscores a statistically significant positive correlation for problems catalogued on GitHub prior to the GPT models’ training cut-off, suggesting that their inclusion in the training data enhanced their performance on related tasks. Concurrent to their work, Li and Flanigan (2023) also plot the performance of several LLMs including GPT-3 (Brown et al., 2020) and open-source LLMs on general natural language understanding tasks against time, specifically before and after each LLM’s training data cutoff date. They find that LLMs are much more likely to improve on a trivial majority baseline for datasets released before their cutoff date, suggesting some contamination. While evaluating on fresh data guarantees a contamination-free performance evaluation, it is hard to scale this method in time as it requires frequent update of evaluation sets. Besides, the work of Cao et al. (2024) shows that evaluating on data posterior to the LLM training cutoff time may not always show a decrease in performance: in the code domain, the opposite may happen, complicating the usage of training cutoffs as a contamination detection method.
Jiang et al. (2024b) delve deep into the impact of evaluation data leakage within the pre-training set by running a series of controlled pre-training experiments with GPT-2 (Radford et al., 2019). They pre-train from scratch GPT-2 in three different ways: with the original pre-training set, with additional input texts from four evaluation datasets (SST-2, MMLU (Hendrycks et al., 2020), CNN/DM (Hermann et al., 2015) and SQuAD-v1 (Rajpurkar et al., 2016)), and with input texts and labels from these same evaluation datasets. Contaminated pre-training clearly boosts performance on all downstream tasks. Findings also reveal a U-shaped relationship between the number of times that contaminated data points are included in the pre-training, and model performance. Authors also show that the n-gram contamination detection method from Llama-2 Touvron et al. (2023) is not effective at detecting contaminated data points, calling for more robust methods.
3.2 Model Memorization
An intuitive approach to detect contamination in LLMs is to frame it as a memorization problem. Researchers focus on which training data points have been memorized verbatim by the model.
Historically, this problem has been addressed in the field of privacy-preserving machine learning: adversarial attacks were crafted as membership inference attacks to access training data points memorized by a black-box language model, potentially compromising sensitive private information. We refer the reader to relevant literature in this research direction (Carlini et al., 2021, 2022).
Elangovan et al. (2021) highlight that some LLMs may benefit from an overlap between training and test sets, and merely memorize data points in this intersection, inflating their test set performance. This overlap is due to the common practice of simple random data shuffling in NLP tasks. By employing a bag-of-words approach to assess text similarity, the study provides a foundational method for identifying and mitigating leakage, though it acknowledges the complexity of semantic similarity measurement and the potential for more sophisticated methodologies to refine overlap detection. Magar and Schwartz (2022) find evidence of contamination across several benchmarks, impacting the performance of models like GPT-3. They propose to quantify the effect of contamination by training on a mix of general and task-specific data, then comparing performance on seen versus unseen instances to measure memorization and exploitation. This methodology highlights the nuanced relationship between memorization and exploitation. Gemma Team Google DeepMind (2024) emphasize the potential vulnerabilities of aligned models to adversarial attacks inducing recitation of memorized training data. Through comparable methods to Anil et al. (2023), they analyze discoverable memorization across 10k documents from the Gemma pre-trained models’ corpora. The analysis distinguishes between exact memorization — where the model’s output precisely matches the source — and approximates memorization, gauged by a 10% edit distance threshold. Gemma models exhibit memorization rates comparable to those of the PaLM and PaLM-2 models.
Recently, researchers have engineered elaborated prompting techniques that elicit data point completion by the LLM, where an output suspiciously close to the actual training data point indicates memorization, which in turn means contamination. Standard prompting asking the LLM to complete a data point may fail due to the LLM’s alignment during RLHF (Ouyang et al., 2022b).
By prompting with an off-(training)-distribution input, Nasr et al. (2023) show that ChatGPT can regurgitate entire chunks of training data, including sensitive personal-level information. The study illustrates that even models trained with alignment techniques, aimed at reducing the emission of memorized data, can be induced to divulge substantial amounts of sensitive information, calling for the development of more robust defenses. The implications of this research are far-reaching, emphasizing the importance of addressing the privacy vulnerabilities inherent in the deployment of LLMs. Weller et al. (2023) highlight the potential for steering LLMs towards generating more factual and grounded content through effective prompting strategies. By appending instructional phrases such as Accoding to that encourage quoting from specific corpora, this study observes improvements in grounding as measured by the Quoted Information Precision (QUIP) Score. This method shows promise across various domains and corpora, indicating its versatility and effectiveness in leveraging LLMs’ memorized knowledge for generating more accurate and reliable responses.
Cloze tasks constitute a special type of prompting where parts of the input is masked. Chang et al. (2023) design a cloze task in which the LLM is prompted with a book’s passage, from which the character name has been replaced by a [MASK] token. Correctly predicting the character’s name out of several passages per book assesses whether the LLM has seen this book during pre-training. The name cloze accuracy is very high for GPT-4 on several popular books such as Alice’s Adventures in Wonderland, and shows a striking difference with BERT’s performance on the same task. Deng et al. (2023) present the TS-Guessing method: one of the options in question-answering benchmarks test sets is hidden, and the LLM is asked to predict it. Exact match rates above 50% for OpenAI models suggest contamination on the MMLU dataset (Hendrycks et al., 2020). The work of Ranaldi et al. (2024) also uses a cloze task: they prompt the LLM to reconstruct columns names in SQL dumps, and compare the performance on an older and more recent dataset to assess contamination. GPT-3.5 accuracy falls from 33.42% to 13.21%, indicating contamination on the older Spider dataset.
3.3 Model Confidence
Going beyond simple string recitation of memorized data points, researchers may analyze the LLM’s confidence to detect contamination: a (too) high confidence may signify prior exposure to the data. Confidence analysis yields several powerful contamination detection techniques, which we highlight require white-box access to the LLM, as researchers need to obtain the output probability distribution for each token during inference.
Next-token likelihood and perplexity (by extension to a whole dataset) are widely used in contamination detection. Given a test sample $X$, the Min-KProb technique (Shi et al., 2024) consists in running the LLM through all tokens of $X$, and keeping track of the $k$tokens with the smallest predicted probability. Then, the average between such bottom probabilities is computed and $X$ is deemed contaminated if this average is too high. Li (2023a) also work on token-level probabilities and compare perplexity on benchmark samples against memorized and clean baselines. The study finds significant memorization in recent models on popular reading comprehension and summarization benchmarks, while multiple-choice benchmarks show less evidence of contamination. This method provides a tool for the community to conduct rigorous contamination analysis, enabling more accurate and reliable model evaluation. Dong et al. (2024) propose two novel likelihood-based contamination detection methodologies: CDD (Contamination Detection via output Distribution) and TED (Trustworthy Evaluation via output Distribution). CDD detects data contamination by observing the peakedness in the LLM’s output distribution in a black-box manner. It represents a significant improvement over existing approaches, offering average relative improvements of 21.8%-30.2% in terms of Accuracy, F1 Score, and AUC metrics. TED corrects the LLM’s output distribution to mitigate the impact of data contamination on evaluation metrics, significantly reducing performance improvements attributed to data contamination across various scenarios and contamination degrees. The work of Xu et al. (2024b) proposes to use perplexity and accuracy on next n-gram prediction. They compute the relative loss in performance on these metrics on paraphrased versions of the dataset created by ChatGPT compared to the original dataset. The difference of relative loss between training and test set is later used to assess contamination or benchmark leakage. Experiments on GSM8K (Cobbe et al., 2021) hint at contamination for QWen (Bai et al., 2023), Aquila and InternLM (Cai et al., 2024) series of LLMs.
Oren et al. (2023) take an original approach to analyze the LLM’s likelihood on an evaluation dataset $D_{E}$. They run inference on $D_{E}$ and on shuffled versions of $D_{E}$. LLM log probabilities being statistically different on the non-shuffled dataset compared to the shuffled versions signifies contamination. The method relies on the assumption that if they are included in the pre-training set, evaluation datasets tend to be present with the same default ordering.
3.4 Model-based Closed-Data Contamination Detection
Relying on an external, high-performing LLM like GPT-4 can be useful in closed-data contamination detection, similarly to the open-data setup.
Golchin and Surdeanu (2023b) make use of GPT-4’s abilities in guided prompting, an enhanced prompting process where the completion prompt includes extra information such as the dataset name. Contamination is then assessed based on the averaged difference in performance between standard and guided prompting, or if GPT-4 with in-context learning finds an exact match or two near-exact matches in the guided completions. This latter method with GPT-4 is highly accurate in identifying contamination (92%–100% accuracy). Moreover, the investigation highlights the prevalence of contamination in datasets such as AG News (Zhang et al., 2015), WNLI (Levesque et al., 2012), and XSum (Narayan et al., 2018), emphasizing the critical need to address data integrity in LLM applications. The same authors follow up with the Data Contamination Quiz (DCQ) evaluation framework for contamination detection in black-box LLMs (Golchin and Surdeanu, 2023a). LLMs are prompted with five completion options, where one option is the ground-truth text from the original dataset, three other options are paraphrases from GPT-4, and the last choice is None. The assumption is that if the LLM picks the exact answer, it is doing so out of memorization, and the authors show that this DCQ framework finds contamination cases than the guided prompting method.
Through the use case of fine-tuning LLMs on mathematical reasoning datasets like GSM8K (Cobbe et al., 2021), Tu et al. (2024) introduce a new paradigm in (closed-data) contamination detection. Their DICE approach consists in finding the layer of the fine-tuned LLM furthest away from its counterpart layer in the not fine-tuned same LLM. Then, a MLP binary classifier is inferred on the isolated layer’s weights to predict contamination. This method is particularly effective for in-distribution contamination detection, a setup defined by the authors as fine-tuning on a dataset from the same distribution as the evaluation dataset. In this setup, DICE reaches near-perfect AUC, outperforming all other methods, including notably Min-KProb.
4 Discussion
4.1 Best Practices
Beyond detecting contaminated evaluation data points, recent works have called the community to adopt better practices to reduce contamination.
Encrypting evaluation datasets
Jacovi et al. (2023) advocate for the encryption of publicly released test data using a public key, coupled with licensing agreements that prohibit derivative distribution. They recommend implementing training exclusion controls for closed APIs and safeguarding test data by refusing to perform evaluations without such controls. Moreover, they suggest avoiding data that appears with the solution and releasing internet-derived data along with the corresponding web-page context.
Scanning newly released evaluation datasets
Sainz et al. (2023) argue to develop automatic or semi-automatic approaches to detect contamination for new benchmarks and design mechanisms to flag those works with contamination.
Evaluating on a wide spectrum
To comprehensively and fairly evaluate LLMs, Zhou et al. (2023) suggest to use more benchmarks from diverse sources, covering both basic and advanced capabilities. Furthermore, they recommend employing multiple task prompts for benchmark tests to derive model performance that is more stable and reliable. Additionally, they highlight the necessity of conducting data decontamination checks between pre-training data and any evaluation data when using benchmarks.
Not leaking data to closed-source APIs
Aside from technical methodologies, Balloccu et al. (2024) conduct a systematic literature review of 225 papers and carefully detail data leakage from them to closed-source models like the GPT-series. Overall, they conclude that $\sim$42of reviewed papers leaked data to GPT-3.5 and GPT-4 for a total of $\sim$4.7M benchmark samples across 263 benchmarks. The authors report evaluation malpractices and propose a list of suggested practices for evaluating closed-source LLMs.
Several data contamination mitigation techniques are detailed in the concurrent survey work of Xu et al. (2024a).
4.2 New Evaluation Benchmarks
Several datasets have been introduced aiming to provide contamination-free evaluation.
Dynamic Evaluation
Some new datasets discard the static paradigm of existing LLM evaluation datasets, and instead propose a dynamic benchmark to avoid contamination:
  • LatestEval (Li et al., 2023b) leverages the most recent texts to create dynamic reading comprehension evaluations. The benchmark is constructed in three steps: (i) collect the latest texts; (ii) extract key information, and (iii) construct questions based on the extracted information through template-filling or LLMs.
  • WIKIMIA (Shi et al., 2024) is a dynamic benchmark made of Wikipedia events created after model training (i.e., after 2023). ChatGPT is leveraged to paraphrase examples for evaluation, as also performed in CleanEval (Zhu et al., 2023).
  • KIEval (Yu et al., 2024) is an interactive evaluation framework incorporating an LLM-powered "interactor", which can ask follow-up questions in multi-round dialogue that leads to contamination-resilient evaluation.
  • LiveCodeBench (Jain et al., 2024) continuously collects new problems from LeetCode, AtCoder and CodeForces to provide a contamination-free code generation benchmark.
Protected Evaluation
Another strategy to avoid contamination of evaluation data is to protect it by locking it out of the public domain. Termite (Ranaldi et al., 2024) is a new text-to-SQL dataset, locked out of public access through search engines via an encryption key. Termite is made of handcrafted databases, each paired with around five queries, designed to match properties of the Spider dataset (Yu et al., 2018), which shows high signs of contamination on GPT-3.5.
Alternative Evaluation
New approaches to LLM evaluation are emerging in the era of widespread public benchmark contamination. The DCQ benchmark utilizes GPT-4 to paraphrase data points from famous benchmarks such as WNLI (Levesque et al., 2012), and checks whether the LLM picks the exact continuation or a paraphrased version, where the former strongly suggests contamination. Moreover, Alzahrani et al. (2024) point out that slight perturbations to existing benchmarks may disrupt model rankings on leaderboards, calling for better evaluation benchmarks.
4.3 Future Challenges
Given the fast and ever-changing landscape of machine learning research, future directions for contamination detection in LLMs could encompass a broad variety of methodologies and technological aspects. We highlight a few potentially critical areas of focus:
Real-time contamination detection systems
Real-time data contamination detection systems that can monitor data streams continuously and alert users to potential contamination events as they happen would be of upmost importance in critical applications like finance (Frésard et al., 2011) and healthcare (Gao et al., 2022) where data integrity is paramount and model reliability critical. Given the sheer volume of new data uploaded daily on the Internet, solving this task requires major technological breakthroughs.
Bypassing contamination detection
Dekoninck et al. (2024) show a very effective way to bypass several existing contamination detection methods: Evasive Augmentation Learning (EAL) consists in paraphrasing benchmarks with GPT-4 and fine-tuning the LLM on the paraphrased data. This method can bypass even oracle-access contamination detection methods, calling for the development of more robust methods.
Ethical and legal data frameworks
We believe in the need for comprehensive ethical and legal frameworks (Chang, 2021) to govern the collection, usage, and management of data for LLM training. This includes policies and protocols for data privacy, consent, and use that help prevent the incorporation of contaminated data from unethical sources and the contamination of widely used pre-training data sources (e.g., CommonCrawl). Developing contamination detection techniques that do not compromise individual privacy will be crucial (Lin et al., 2010; Hayes and Ohrimenko, 2018).
5 The LLMSanitize Library
LLM ARC HellaSwag MMLU TruthulQA Winogrande GSM8K
Guided Prompting (ROUGE-L)
Llama-2-7B 0.064 0.489 0.146 0.001 0.001 0.968
Qwen-1.5-7B 0.000 0.023 0.700 0.664 0.202 1.000
Mistral-7B 0.000 0.000 0.010 0.042 0.999 0.444
Gemma-7B 0.978 0.000 0.580 0.122 0.820 0.000
GPT-3.5-turbo 0.000 0.184 0.970 0.075 0.000 1.000
Sharded Likelihood
Llama-2-7B 0.258 0.000 0.967 0.154 0.171 0.781
Qwen-1.5-7B 0.426 0.000 0.476 0.255 0.115 0.209
Mistral-7B 0.862 0.000 0.376 0.057 0.485 0.507
Gemma-7B 0.234 0.000 0.255 0.450 0.215 0.648
Table 1: P-value for two closed-data contamination detection methods using a statistical test: Guided Prompting (Golchin and Surdeanu, 2023a) (top) and Sharded Likelihood (Oren et al., 2023) (bottom). Shaded results are not statistically significant. We ran LLMSanitize on 100 test data points.
To facilitate progress in contamination detection in LLMs, we have built and are releasing an open-source Python library, named LLMSanitize. LLMSanitize supports most of the open-data and closed-data contamination detection described above and listed in Figure 1. It relies on the HuggingFace hub to access LLMs and the data, namely transformers (Wolf et al., 2020) and datasets (Lhoest et al., 2021) libraries. Efficient LLM inference for model-based use cases is handled with the vLLM library (Kwon et al., 2023). LLMSanitize is straightforward to install through PyPI. Users are expected to pass in the relevant arguments for each contamination detection method, as well as key parameters such as the LLM and dataset name and the number of data points to run the contamination detection on. We refer to Appendix A.
We now put to test LLMSanitize. We apply Guided Prompting (Golchin and Surdeanu, 2023b) and Sharded Likelihood (Oren et al., 2023) on the six datasets of HuggingFace OpenLLM Leaderboard [^1] https://huggingface.co/open-llm-leaderboard. We employ four widely used 7B-size LLMs: Llama-2 (Touvron et al., 2023), Qwen-1.5 (Bai et al., 2023), Mistral (Jiang et al., 2023) and Gemma (Gemma Team Google DeepMind, 2024). We also evaluate GPT-3.5-turbo on guided prompting - the blocked access to the logprobs prohibits us from applying the other method. Results are shown in Table 1. The HellaSwag benchmark (Zellers et al., 2019) shows strong signs of contamination, consistent across methods and LLMs.
6 Conclusion
As LLMs rapidly evolve and the training data become more extensive, LLMs’ performances are inevitably biased by contamination. Therefore, in this survey, we investigated and structured the current landscape of contamination detection in LLMs. We introduced a new Python library named LLMSanitize which is a universal tool for quickly detecting contamination in various models and datasets. The extensive investigation and library could provide a universal foundation for future models when considering contamination issues.
Limitations
In this paper, we reviewed all work related to the detection of contamination. Contamination itself, as a general problem with LLMs, is larger than detection. For instance, our study scope does not focus on the mitigation of contamination, which we only briefly touch in Section 4.1. Mitigation is also of critical importance ; but we made the choice in this study to narrow down and focus on detection, which we believe makes our contribution more focused.
We included all works mentioning the issue of detecting or assessing contamination with language models from the last four years. Many of these works, for instance language models reports (Radford et al., 2019), are not focusing on the issue of contamination detection: these papers simply discuss this issue in a few paragraphs. Our survey also includes papers entirely dedicated to contamination detection, such as the LLM Decontaminator (Yang et al., 2023b). The technical depth of the work related to contamination detection across such differing types of works is very heterogeneous. This is not reflected in our survey as we chose to include all papers on the same level.
Appendix A LLMSanitize Usage
In this section we showcase how to use our Python library LLMSanitize. The library is recommended to be used with Python 3.9 and CUDA 11.8. The first step is to create the environment and install the library.
The following code should be run in a terminal:
1conda create --name llmsanitize python=3.9
2conda activate llmsanitize
3pip install llmsanitize==0.0.8
We first review how to use LLMSanitize for an open-data contamination detection setup. In this case, we compare two datasets and no LLM is involved. Let us take a simple setup: we compare HellaSwag as evaluation dataset against ARC as training dataset, sub-sampling 100 evaluation data points, and using GPT-2’s 8-grams collision as contamination detection method.
The following Python code could be run:
1import argparse
2from llmsanitize import OpenDataContaminationChecker
3
4parser = argparse.ArgumentParser()
5parser.add_argument("--train_data_name", type=str, default="allenai/ai2_arc")
6parser.add_argument("--eval_data_name", type=str, default="Rowan/hellaswag")
7parser.add_argument("--eval_set_key", type=str, default="validation")
8parser.add_argument("--text_key", type=str, default="ctx")
9parser.add_argument("--n_eval_data_points", type=int, default=100)
10parser.add_argument("--method", type=str, default="gpt-2")
11args = parser.parse_args()
12
13contamination_checker = OpenDataContaminationChecker(args)
14contamination_checker.run_contamination(args.method)
From there, the library will run the contamination detection and log results to a dedicated file. We refer the reader to all argument fields defined in the main.py file ; which among other properties, provide an option to stream over the (large) training dataset in order to limit memory usage for the use case of open-data contamination.
Next, we review how to conduct a closed-data contamination detection use case, involving an LLM and an evaluation dataset. For this setup, we apply the Guided Prompting method on Llama-2 on the ARC dataset. For this method (and all closed-data contamination detection methods except Sharded Likelihood), we make use of vLLM for efficient inference.
This requires us to first open a vLLM instance. Run the following shell script, which assumes access to 2 GPUs:
1export LD_LIBRARY_PATH=/usr/local/cuda/lib64:\$LD_LIBRARY_PATH
2export RAY_memory_monitor_refresh_ms=0;
3export CUDA_VISIBLE_DEVICES=0,1;
4server_type=vllm.entrypoints.openai.api_server
5
6python -m \$server_type \
7 --model <path_to_local_llama2> \
8 --tokenizer <path_to_local_llama2> \
9 --gpu-memory-utilization=0.9 \
10 --max-num-seqs=200 \
11 --disable-log-requests \
12 --host 127.0.0.1 \
13 --port <port_number> \
14 --tensor-parallel-size 2 \
15 --download-dir /export/home/cache
Next, a minimal Python code to run Guided Prompting detection could be:
1import argparse
2from llmsanitize import ClosedDataContaminationChecker
3
4parser = argparse.ArgumentParser()
5parser.add_argument("--eval_data_name", type=str, default="allenai/ai2_arc")
6parser.add_argument("--eval_set_key", type=str, default="test")
7parser.add_argument("--text_key", type=str, default="question")
8parser.add_argument("--label_key", type=str, default="anskwerKey")
9parser.add_argument("--n_eval_data_points", type=int, default=100)
10parser.add_argument("--num_proc", type=int, default=100)
11parser.add_argument("--method", type=str, default="guided-prompting")
12parser.add_argument("--local_port", type=int, default=<port_number>)
13parser.add_argument("--model_name", type=str, default=<path_to_local_llama2>)
14parser.add_argument("--guided_prompting_task_type", type=str, default="CLS")
15parser.add_argument("--user_local_model", type=bool, default=True)
16args = parser.parse_args()
17
18contamination_checker = ClosedDataContaminationChecker(args)
19contamination_checker.run_contamination(args.method)
Note that the model path and port number passed as arguments in the Python script should match the ones used to launch the vLLM instance.